# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v1

@@ 3 WaDiRo SCNNs
This is a common assumption in statistical machine learning which is also made by, e.g., [36, 25]. Let $d_{ji}=\mathbf{D}_{i}(j,j),\mathbf{d}_{j}\in\{0,1\}^{P}$. We consider the modified features ${\hat{\mathbf{x}}}_{j}^{\top}=\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top}\right)\ \forall j\in\llbracket N\rrbracket$.
@@
Suppose that the training loss function $\mathcal{L}$ is the $\ell_{1}$-norm and let $\hat{\mathbf{z}}_{j}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)\ j\in\llbracket N\rrbracket$ be modified samples. Then, the convex training procedure (SCNNT) is equivalent to a linear regression problem.
@@
$\mathcal{L}\left(\sum_{i=1}^{P}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i}),\mathbf{y}\right)$ $=\left\|\left(\sum_{i=1}^{P}d_{ji}\mathbf{x}_{j}^{\top}\mathbf{u}_{i}\middle)\right|_{j=1}^{N},\mathbf{y}\right\|_{1}$
@@
$=\sum_{j=1}^{N}\left|\left(\operatorname{vec}(\mathbf{U})^{\top},-1\right)\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},y_{j}\right)^{\top}\right|$
$=\sum_{j=1}^{N}\left|\left(\operatorname{vec}(\bm{\nu})^{\top},\operatorname{vec}(-\bm{\omega})^{\top},-1\right)\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},y_{j}\right)^{\top}\right|$
@@
where $\bm{\beta}=\left(\operatorname{vec}(\bm{\nu})^{\top},\operatorname{vec}(-\bm{\omega})^{\top},-1\right)$ represent the weights of the linear regression, i.e., the optimization variable of the learning problem and ${\hat{\mathbf{z}}}_{j}=\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},y_{j}\right)^{\top}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)^{\top}\ \forall j\in\llbracket N\rrbracket$ are the modified samples. The training problem (SCNNT) can be written as:
$p^{\star}_{\ell_{1}}=$ $\min_{\bm{\nu},\bm{\omega}:\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\ \forall i\in\llbracket P\rrbracket}$ $\left\|\mathbf{y}-\sum_{D_{i}\in\mathcal{D}_{\mathbf{X}}}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i})\right\|_{1}$
$=$ $\min_{\bm{\nu},\bm{\omega}:\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\ \forall i\in\llbracket P\rrbracket}$ $\sum_{j=1}^{N}\left|{\bm{\beta}}^{\top}{\hat{\mathbf{z}}}_{j}\right|,$
and is equivalent to a linear regression problem with respect to the modified samples. $\square$
The absolute error loss function tends to be more robust against outliers when training regression models [36]. It is then a reasonable choice in an application where distributional robustness is desired and will be used for the remainder of this work.
We remark that the transformed feature space preserves the independence of individual samples when conditioned on $\mathcal{S}=\{\mathbf{S}_{1}=\mathbf{s}_{1},\mathbf{S}_{2}=\mathbf{s}_{s},\ldots,\mathbf{S}_{P}=\mathbf{s}_{P}\}$. Whether $\mathcal{S}$ is constructed by random sampling or by deterministic hyperplane arrangements has no importance as long as these vectors are saved prior to training and are used throughout the training and evaluation.
@@
Suppose Assumption 1 holds, then $\hat{\mathbf{z}}|\mathcal{S}$ is independent and identically distributed.
@@
$\mathbf{x}_{j}\mathbf{d}_{j}^{\top}$ $=\mathbf{x}_{j}\begin{bmatrix}\mathbbm{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{1}\geq 0)&\mathbbm{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\mathbbm{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}\quad$ $\forall j\in\llbracket N\rrbracket.$
@@
$g_{\bm{\beta}}(\hat{\mathbf{z}}_{j})$ $=\left|\bm{\beta}^{\top}\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},y_{j}\right)^{\top}\right|$ $\forall j\in\llbracket N\rrbracket$
$=\left|\bm{\beta}^{\top}\left(f(\mathbf{x}_{j},\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{P}),y_{j}\right)^{\top}\right|$ $\forall j\in\llbracket N\rrbracket,$
where $\hat{\mathbf{x}}_{j}=f(\mathbf{x}_{j},\mathcal{S})$. We have shown that each $\hat{\mathbf{x}}_{j}$ is solely function of $\mathbf{x}_{j}$ and $\mathcal{S}$. Because $\mathcal{S}$ is obtained prior to training and the mapping $f$ is used similarly on every sample, $\hat{\mathbf{z}}|\mathcal{S}$ is independent and identically distributed.  $\square$
Following Lemmas 1 and 2, we consider the new empirical distribution of the modified samples conditioned on $\mathcal{S}$: $\hat{\mathbb{F}}_{N}^{\mathcal{S}}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\hat{\mathbf{z}}_{j}|\mathcal{S}}(\hat{\mathbf{z}}|\mathcal{S})$, where $\hat{\mathbf{z}}_{j}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)^{\top}$. We formulate the non-tractable Wasserstein distributionally robust training problem of the SCNN by centering the Wasserstein ball on $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$:
@@
where $\hat{\mathbf{z}}\in\hat{\mathcal{Z}}$ and $\mathcal{P}(\hat{\mathcal{Z}})$ is the space of all possible distributions with support on $\hat{\mathcal{Z}}=\mathbb{R}^{2Pd+1}$.
@@
    s.t.
    ${\bm{\beta}}_{i}$ $\leq a\quad$ $\forall i\in\llbracket 2Pd+1\rrbracket$
    $-{\bm{\beta}}_{i}$ $\leq a\quad$ $\forall i\in\llbracket 2Pd+1\rrbracket$
@@
    s.t.
    $\|{\bm{\beta}}\|_{2}^{2}$ $\leq a^{2}\quad\quad$
@@
Proof: Lemma 1 states that the training problem, with respect to the modified samples $\hat{\mathbf{z}}$, is similar to a constrained linear regression. Lemma 2 shows that the distribution of the modified samples is i.i.d. when conditioning on $\mathcal{S}$. Because (WDR-SCNNT) uses $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$, by invoking Lemmas 1 and 2, we reformulate (WDR-SCNNT) as a convex program in the form of (TWDRO). The final forms then follows from [25, Theorem 6.3] and [36, Chapter 4]. The detailed proof is provided in Appendix A for completeness. $\square$
We remark that (WDR1-SCNNT) and (WDR2-SCNNT) can be easily modified to take into account the bias weights of the hidden and output layers of the SNN. The process is as follows: (i) extend $\mathbf{X}$ with a column of ones which is equivalent to adding bias weights in the hidden layer; (ii) let $\bm{\beta}$ be defined as $\left(\operatorname{vec}(\bm{\nu})^{\top},\operatorname{vec}(-\bm{\omega})^{\top},b,-1\right)$ and $\hat{\mathbf{z}}_{j}$ as $\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})^{\top},1,y_{j}\right)^{\top}\ \forall j\in\llbracket N\rrbracket$ where $b\in\mathbb{R}$ is the bias of the sole output neuron. We also remark that both formulations of Proposition 1 are fairly fast to solve with open-source solvers, e.g., Clarabel [39]. By using these formulations to train SNNs, we efficiently obtain an optimal distributionally robust shallow neural network. Recall that the weights from the standard non-convex SNN training problem can be retrieved via (2) and (3).

### v2

@@ 3 Wasserstein Distributionally Robust Shallow Convex Neural Networks
This is a common assumption in statistical machine learning which is also made by, e.g., Chen and Paschalidis (2020) and Mohajerin Esfahani and Kuhn (2018). In practice, in learning settings, we only have access to a finite amount of training samples from which we can indirectly observe $\mathbb{P}$. This motivates the use of a finite empirical distribution ^$\hat{\mathbb{P}}_{N}$ that approximates $\mathbb{P}$ (Kuhn et al. 2019). As such, we assume that our empirical samples are drawn independently from a consistent unknown distribution which is i.i.d. When formulating (TWDRO), only the independence of the samples is necessary to build ^$\hat{\mathbb{P}}_{N}$. The stronger identically distributed assumption is then only required to bound the out-of-sample performance, see Section 4. As a result, the later part of Assumption 1 implies that any distribution drift is neglected from the analysis. Recall (1) and let $d_{ji}=\mathbf{D}_{i}(j,j)=\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{i}\geq\mathbf{0}),\ \forall j\in\llbracket N\rrbracket,\ i\in\llbracket P\rrbracket$ such that $\mathbf{d}_{j}\in\{0,1\}^{P}$ is the vector encoding the activation state of $\mathbf{x}_{j}$ for each vector in $\mathcal{S}$. For the remainder of this section, we consider the modified features ${\hat{\mathbf{x}}}_{j}=\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})\right)\ \forall j\in\llbracket N\rrbracket$, which are a higher order basis expansion of the original features with their respective possible activation patterns.
We now proceed to decouple the training procedure for distinct samples.
@@
Suppose that the training loss function $\mathcal{L}$ is the $\ell_{1}$-norm, and let $\hat{\mathbf{z}}_{j}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)\ j\in\llbracket N\rrbracket$ be modified samples. Then, the convex training procedure (SCNNT) is equivalent to a constrained linear regression problem.
@@
$\mathcal{L}\left(\sum_{i=1}^{P}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i}),\mathbf{y}\right)$ $={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\|\left(\sum_{i=1}^{P}d_{1i}\mathbf{x}_{1}^{\top}\mathbf{u}_{i},\sum_{i=1}^{P}d_{2i}\mathbf{x}_{2}^{\top}\mathbf{u}_{i},\ldots,\sum_{i=1}^{P}d_{Ni}\mathbf{x}_{N}^{\top}\mathbf{u}_{i}\right)-\mathbf{y}\right\|_{1}}}$
@@
$=\sum_{j=1}^{N}\left|\left(\operatorname{vec}(\mathbf{U}),-1\right)^{\top}\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),y_{j}\right)\right|$
$=\sum_{j=1}^{N}\left|\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)^{\top}\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),y_{j}\right)\right|$
@@
where $\bm{\beta}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)$ represent the weights of the linear regression and ${\hat{\mathbf{z}}}_{j}=\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),y_{j}\right)=\left(\hat{\mathbf{x}}_{j},y_{j}\right)\ \forall j\in\llbracket N\rrbracket$ are the modified samples. The training problem (SCNNT) can be written as:
$p^{\star}_{\ell_{1}}=$ $\min_{\bm{\nu},\bm{\omega}:\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\ \forall i\in\llbracket P\rrbracket}$ $\left\|\mathbf{y}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{i=1}^{P}}}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i})\right\|_{1}$
$=$ $\min_{\bm{\nu},\bm{\omega}:\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\ \forall i\in\llbracket P\rrbracket}$ $\sum_{j=1}^{N}\left|{\bm{\beta}}^{\top}{\hat{\mathbf{z}}}_{j}\right|$
s.t. $\bm{\beta}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right),$
and is equivalent to a constrained linear regression problem with respect to the modified samples. Additionally, we note that the LASSO and Ridge regularized SCNN training problems are exactly equivalent to their respective linear regression formulation.
$p^{\star}_{\ell_{1},r}=$ $\min_{\bm{\nu},\bm{\omega}:\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\ \forall i\in\llbracket P\rrbracket}$ $\left\|\mathbf{y}-\sum_{D_{i}\in\mathcal{D}_{\mathbf{X}}}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i})\right\|_{1}+\sum_{i=1}^{P}\left(r(\bm{\nu}_{i})+r(\bm{\omega}_{i})\right)$
$=$ $\min_{\bm{\nu},\bm{\omega}:\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\ \forall i\in\llbracket P\rrbracket}$ $\sum_{j=1}^{N}\left|y_{j}-\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega})\right)^{\top}{\hat{\mathbf{x}}}_{j}\right|+r\left([\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega})]\right),$
where $r(\cdot)=\|\cdot\|_{1}$ for LASSO regularization and $r(\cdot)=\|\cdot\|_{2}^{2}$ for Ridge regularization. $\square$
The $\ell_{1}$-norm loss function tends to be more robust against outliers than other regression objectives when training regression models (Chen and Paschalidis 2020). It is then a preferable choice in an application where distributional robustness is desired and will be used for the remainder of this work as it is also convenient in the WaDiRo reformulation. Nonetheless, we remark that Lemma 1 can be adapted to other separable loss functions, i.e., functions $f$ such that $f(\phi_{1}(x_{1})+\phi_{2}(x_{2})+\ldots\phi_{N}(x_{N}))=f(\phi_{1}(x_{1}))+f(\phi_{2}(x_{2}))+\ldots+f(\phi_{N}(x_{N}))$ where $x_{j}\ j\in\llbracket N\rrbracket$ are variables and $\phi_{j}\ j\in\llbracket N\rrbracket$ are arbitrary functions.
We remark that the transformed feature space preserves the independence of individual samples when conditioned on $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{P}\}$ because the mapping from the original feature space to the modified one depends only on individual samples and $\mathcal{S}$. Whether $\mathcal{S}$ is constructed by random sampling or a by deterministic hyperplane arrangement has no importance as long as these vectors are saved prior to training and are used throughout the training and evaluation. From this observation, we now introduce our second lemma.
@@
Suppose Assumption 1 holds, i.e., samples $\mathbf{z}\in\mathcal{Z}$ are i.i.d., then modified samples given the set $\mathcal{S}$, $\hat{\mathbf{z}}|\mathcal{S}\in\hat{\mathcal{Z}}$, are also i.i.d.
@@
$\mathbf{x}_{j}\mathbf{d}_{j}^{\top}$ $=\mathbf{x}_{j}\begin{bmatrix}\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}\quad$ $\forall j\in\llbracket N\rrbracket.$
@@
$g_{\bm{\beta}}(\hat{\mathbf{z}}_{j})$ $=\left|\bm{\beta}^{\top}\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),y_{j}\right)\right|$ $\forall j\in\llbracket N\rrbracket$
$=\left|\bm{\beta}^{\top}\left(f(\mathbf{x}_{j},\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{P}),y_{j}\right)\right|$ $\forall j\in\llbracket N\rrbracket$
$=\left|\bm{\beta}^{\top}\left(\hat{\mathbf{x}}_{j},y_{j}\right)\right|$ $\forall j\in\llbracket N\rrbracket,$
where $\hat{\mathbf{x}}_{j}=f(\mathbf{x}_{j},\mathcal{S})$ is a basis expansion of the original feature space using the definition of the hyperplane generation (or approximation) and $g_{\bm{\beta}}(\hat{\mathbf{z}}):\hat{\mathcal{Z}}\times\mathbb{R}^{2Pd+1}\to\mathbb{R}$ is the modified objective function when applying $\bm{\beta}$ on a modified sample. We have showed that each $\hat{\mathbf{x}}_{j}$ depends solely on $\mathbf{x}_{j}$ and $\mathcal{S}$. Because $\mathcal{S}$ is obtained prior to training and the mapping $f$ is the same on every sample, $\hat{\mathbf{z}}|\mathcal{S}$ is also i.i.d.  $\square$
Following Lemmas 1 and 2, we consider the new empirical distribution of the modified samples conditioned on $\mathcal{S}$: $\hat{\mathbb{F}}_{N}^{\mathcal{S}}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\hat{\mathbf{z}}_{j}|\mathcal{S}}(\hat{\mathbf{z}}|\mathcal{S})$, where $\hat{\mathbf{z}}_{j}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)^{\top}$. As the number of modified samples grows, $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$ converges to the true underlying distribution $\mathbb{F}^{\mathcal{S}}$.
To make our case for the similarity between samples of ^$\hat{\mathbb{P}}_{N}$ and $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$, we provide a detailed example of the mapping between the samples of the two distributions.
Example 1.
Let $\mathbf{z}_{j}=(\mathbf{x}_{j},y_{j})$ be any sample from ^$\hat{\mathbb{P}}_{N}$ and $\hat{\mathbf{z}}_{j}=(\hat{\mathbf{x}}_{j},y_{j})$ be the equivalent sample from $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$. Suppose that $\operatorname{card}(\mathcal{S})=3$ and that $\mathbf{d}_{j}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{2}\geq 0)&\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{3}\geq 0)\end{bmatrix}=\begin{bmatrix}1&0&1\end{bmatrix}$, then:
[…]
and accordingly:
[…]
Now, recall that $\mathcal{S}$ is known a priori and can be generated randomly to approximate the hyperplane arrangement of ^$\hat{\mathbb{P}}_{N}$. As such, $\mathcal{S}$ do not bring additional information to $\mathbf{x}_{j}$; indeed $\hat{\mathbf{x}}_{j}$ is purely an expansion of $\mathbf{x}_{j}$ with a possible $\operatorname{ReLU}$ activation pattern. Effectively, we map the original samples to a higher dimension and apply the Wasserstein distributionally robust regularization on this higher representation which is closely related to the original feature space.
We formulate the non-tractable Wasserstein distributionally robust training problem of the SCNN by centering the Wasserstein ball on $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$:
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@
where $\hat{\mathbf{z}}\in\hat{\mathcal{Z}}$ and $\mathcal{P}(\hat{\mathcal{Z}})$ is the space of all possible distributions with support on $\hat{\mathcal{Z}}=\mathbb{R}^{2Pd+1}$. Note that $g_{\bm{\beta}}(\hat{\mathbf{z}})=\left|\bm{\beta}^{\top}\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),y_{j}\right)\right|$ as defined in Lemma 2.
@@
    s.t. ${\bm{\beta}}_{k}$ $\leq a\quad$ $\forall k\in\llbracket 2Pd+1\rrbracket$
    $-{\bm{\beta}}_{k}$ $\leq a\quad$ $\forall k\in\llbracket 2Pd+1\rrbracket$
@@
    ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}}}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@
    s.t. $\|{\bm{\beta}}\|_{2}^{2}$ $\leq a^{2}\quad\quad$
@@
    ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}}}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@
Proof: Lemma 1 states that the training problem, with respect to the modified samples $\hat{\mathbf{z}}$, is similar to a constrained linear regression over a basis expansion of the original feature space. This problem is convex in its objective and its constraints. Lemma 2 shows that the distribution of the modified samples is i.i.d. when conditioned on $\mathcal{S}$. Because (WDR-SCNNT) centers the Wasserstein ball on $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$, by invoking Lemmas 1 and 2 regarding the modified sample space, we can reexpress (WDR-SCNNT) to the tractable DRO formulation (TWDRO). The final forms then follow from Theorem 6.3 of  Mohajerin Esfahani and Kuhn (2018) and are inherited from WaDiRo linear regressions (Chen and Paschalidis 2020, Chapter 4). The detailed proof, adapted to shallow convex neural networks, is provided in Appendix A for completeness. $\square$
We remark that (WDR1-SCNNT) and (WDR2-SCNNT) can be easily modified to take into account the bias weights of the hidden and output layers of the SNN. The process is as follows: (i) extend $\mathbf{X}$ with a column of ones which is equivalent to adding bias weights in the hidden layer; (ii) let $\bm{\beta}$ be defined as $\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),b,-1\right)$ and $\hat{\mathbf{z}}_{j}$ as $\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),1,y_{j}\right)\ \forall j\in\llbracket N\rrbracket$ where $b\in\mathbb{R}$ is the bias of the sole output neuron. We also remark that both formulations of Proposition 1 are fairly fast to solve with open-source convex solvers, e.g., Clarabel (Goulart and Chen 2024), if the maximal number of neurons dictated by $\tilde{\mathcal{D}}_{X}$ and the size of the training set are reasonable in regards to the available computational power, e.g., in our case less than 300 neurons and 3000 samples were required. We provide time comparisons in Section B.3 and identify acceleration methods from the literature that could be adapted to our formulation in Section 9. Moreover, with only two hyperparameters, i.e., the maximal number of neurons and the Wasserstein ball radius, the hyperparameter search space is significantly reduced compared to standard non-convex training programs. As such, they contribute to a lighter MLOps pipeline by diminishing the computational effort required during hyperparameter tuning. Recall that the weights from the standard non-convex SNN training problem can be retrieved via (2) and (3) even though they are not required to obtain accurate predictions as the SCNN representation is sufficient.